OpenStudy (anonymous):

whats the indefinite integral of sin3x cos 3x dx? Also can I get a link of the identities? my teacher has included 1 and the textbook only 1 or 2. thanks!

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ 1 }{ 2 } \sin 6t +\sin(0)\] which gives \[\frac{ 1 }{ 12 } \cos 6x + c\] But the correct answer is \[\frac{ -1 }{ 12 } \cos 6x + c\] Please explain

OpenStudy (anonymous):

Integral of sin 6x is -cos 6x /6.

OpenStudy (anonymous):

oh right..rookie mistake, btw do u have any link to identies containing cos and sin?